# Probing Mechanical Reasoning in Large Vision Language Models

### v3

@@ Abstract
Mechanical reasoning is a hallmark of human intelligence, defined by its ubiquitous yet irreplaceable role in human activities ranging from routine tasks to civil engineering. Embedding machines with mechanical reasoning is therefore an important step towards building human-level artificial intelligence. Here, we leveraged 155 cognitive experiments to test the understanding of system stability, gears and pulley systems, leverage principle, inertia and motion, and fluid mechanics in 26 Vision Language Models (VLMs). Results indicate that VLMs consistently perform worse than humans on all domains, while demonstrate significant difficulty in reasoning about gear systems and fluid mechanics. Notably, their performance on these tasks do not improve as number of parameters increase, suggesting that current attention-based architecture may fail to grasp certain underlying mechanisms required for mechanical reasoning, particularly those pertaining to mental simulations.
Keywords: mechanical reasoning; vision language models; model-based reasoning; intuitive physics; cognitive AI
@@ 1 Introduction
Humans are uniquely capable of working with complex mechanical systems, ranging from routine tasks, such as assembling furniture, to large-scale civil endeavors, such as designing architectural structures and developing advanced technologies (Harari, 2014). These capabilities are underpinned by the cognitive ability to reason about the relationships and interactions of physical objects, an ability known as mechanical reasoning (Clark, 2010; Harman, 2011; Vaesen, 2012). While there is evidence that some animal species demonstrate limited mechanical reasoning, such as understanding basic tool use or object interactions (Shumaker et al., 2011), the human capacity for mechanical reasoning is unparalleled in its flexibility, sophistication, and creative potential. This distinction allows humans to innovate, solve complex problems, and adapt tools and systems to a wide range of environments and challenges (Allen et al., 2020; Allen, 2021). As such, mechanical reasoning is a cornerstone of human intelligence, driving technological and cultural progress throughout history. Implementing mechanical reasoning is therefore a vital step toward developing artificial intelligence systems capable of achieving human-level performance in real-life scenarios. Given the rapid advancements in large language models (LLMs), particularly their variants that support visual operations, i.e. vision language models (VLMs), mechanical reasoning appears to be a critical area of assessment to evaluate their current reasoning capabilities and identify potential limitations.
Although mechanical reasoning is a high-level cognitive ability that does not emerge until late childhood, it has been found to rely heavily on more foundational cognitive strategies (Hegarty et al., 1988; Kim & Spelke, 1999; Allen, 2021). In particular, decades of research in cognitive science suggest that mental simulation, the process of constructing and operating mental models of the world to guide reasoning, is critical for many aspects of mechanical reasoning (Hegarty, 2004). For example, studies employing methods such as interviews and eye-tracking have revealed that individuals dynamically construct spatial representations of gear and pulley systems to infer their properties as they mentally simulate stages of motion (Hegarty, 1992; Lehrer & Schauble, 1998; Kubricht et al., 2017; Rozenblit et al., 2002). On the other hand, while it remains an ongoing debate, several studies have indicated that the ability to construct mental models of the world to guide reasoning, particularly in the context of spatial dimensions, remains significantly underdeveloped even in state-of-the-art LLMs (Hao et al., 2023; Mitchell & Krakauer, 2023; Goddu et al., 2024; Gao et al., 2024). Since mental simulations necessitate this model-based reasoning capacity on the visuospatial domain, a thorough evaluation of the mechanical reasoning capabilities of VLMs could therefore provide valuable insights into this debate given the critical role of mental simulation in solving mechanical reasoning tasks.
In the present work, to systematically evaluate VLMs’ mechanical reasoning abilities, we constructed the MechBench, which contains around 155 cognitive experiments. The experiments investigate six generalized domains of mechanical reasoning: system ability, pulley systems, gear systems, leverage principle, inertia and motion, and fluid mechanics.
Figure 1: Sample Tasks on the Six Domains of MechBench. For each domain, GPT4o answered correctly to the task on the left and failed the task on the right
@@ 2 Methods
2.1 Cognitive Experiments
Mechanical reasoning has been widely explored in cognitive psychology and educational research. This study uses single-image adaptations of classic cognitive tasks from the literature to assess mechanical reasoning, focusing on six key aspects: system stability, pulley systems, gear systems, leverage principle, inertia and motion, and fluid mechanics. Example cases on each domains are presented in Figure 1. This categorization is intentionally designed to encompass the full range of mechanical reasoning, ensuring each category represents a distinct but fundamental dimension of physical understanding. Below, we provide explanations for each domain.
@@ 2.1 Cognitive Experiments
Fluid Mechanics
Fluid mechanics involves understanding the behavior of liquids under various conditions, such as flow, external pressures, and volume changes. Although gounded in the intuitive understanding of fluid dynamics emerged very early in humans’ cognitive development (Hespos et al., 2016), reasoning about liquid behaviors in mechanical systems require simultaneous consideration of geometry, force, and dynamics. These tasks represent a highly important dimension of mechanical reasoning concerning specifically about liquid motions as opposed to solid objects alone.
2.2 Model Selection and Experiment
@@
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series(Hong et al., 2024), Qwen series(Qwen-VL (Bai et al., 2023), Qwen-2 (Wang et al., 2024)), and Blip2 (Li et al., 2023), LLaVA-Next (Liu et al., 2024) , which are capable of reasoning over interleaved multiple images and texts.
@@ 2.3 Human Baseline
We recruited a total of 9 participants, all of whom were college students proficient in English. Participants were instructed to skip any question that was ambiguously phrased or too complex to answer within 90 seconds. A question was marked as failed if the participant did not provide an answer. For each question, at least 80% of participants needed to answer correctly; otherwise, we modified the question, and new annotators completed the revised version. The human baseline result for each question was normalized based on the number of participants who provided an answer.
In addition to answering the experimental questions, participants completed a test of visuospatial fluid intelligence using Raven’s Advanced Progressive Matrices (APM), a widely used non-verbal assessment tool for evaluating fluid intelligence (Raven et al., 2000). The APM results were collected to explore correlations between participants’ general visuospatial reasoning abilities and their performance on the experimental tasks.
@@ 3 Results
3.1 General Results
Our study reveals a significant disparity between human and model performance across multiple evaluation dimensions. As shown in Figure 2 and 3, humans consistently outperform models both in overall accuracy and in task-specific dimensions. These results highlight the limitations of current VLMs in replicating human-like reasoning in mechanical and intuitive physics tasks. Among the evaluating dimensions, Pulley Systems exhibits the largest performance gap, with human accuracy nearing 90%, compared to model accuracy averaging around 50%.
@@ 3.1 General Results
Figure 3: Overall and Dimension-Wise Accuracy: Humans vs. Models. A. For overall accuracy across tasks, human participants outperform models significantly (p < 0.0001). B. Human participants consistently outperform models in each dimension (all categories p < 0.001).
3.2 Human vs. Model Performance Across Each Dimensions
The ANOVA test ($F=11.8111,p<0.0001$) revealed significant differences in human performance across the six categories. Tukey HSD post-hoc analysis (visualized in the heatmap) shows that the performance of humans handling the pulley system is significantly different from the other five categories. As can be seen from the heat map, the performance of humans handling the pulley system is significantly better than the other five categories. In contrast, ANOVA results for model performance ($F=1.6809,p=0.1430$) indicated no significant differences across task categories. Tukey HSD analysis further reveals that models exhibited similar accuracy across all categories with the exception of Fluid System, which deviates significantly with other dimensions. Together, as shown in Figure 3, the distribution of human performance is narrower, indicating more consistent accuracy, whereas the models display a wider range of variability.
Figure 4: Comparison of Tukey HSD P-values for Model and Human Performance Across Different Dimensions. A. The ANOVA results for model performance reveal a significant difference across the six dimensions (F-statistic = 11.8111, p ¡ 0.0001), indicating that the models exhibit varying accuracy depending on the task. Tukey HSD tests show that significant differences exist between several pairs of dimensions, with System Stability, Inertia and Motion, and Fluid Systems exhibiting large disparities in accuracy. Leverage Principle and Gear Systems show less pronounced differences. B. Human performance across the six dimensions shows no significant difference (F-statistic = 1.6809, p = 0.1430). Tukey HSD tests for human performance indicate that none of the comparisons between different dimensions reach statistical significance, suggesting a more uniform performance across dimensions.
@@ 3.3 Relationship Between Model Performance and Model Size
A widely held belief in the machine learning community is that an increase in a model’s scale, measured by the number of parameters, leads to systematic improvements in its reasoning abilities (Sutton, 2019; Kaplan et al., 2020), a concept known as the scaling law. However, this assumption is an empirical observation without theoretical proof. To evaluate whether the scaling law applies to mechanical reasoning, we further examined the relationship between model performance on mechanical reasoning tasks and model size, as measured by the number of parameters (Figure 5). For overall accuracy across task, regression analysis yielded a formula $y=0.0821x+0.4318$ (p = 0.0053, R² = 0.2917). This trend was particularly evident in models with parameter counts exceeding 10 billion, such as GPT-4o High and GPT4V-TURBO-HIGH, which consistently outperformed smaller models. In most of the dimensions (System Stability, Pulley Systems, Leverage Principle, Inertia and Motion, Fluid Mechanics), accuracy increases with larger model sizes. However, Gear Systems and Fluid Systems show very weak relationships, with slopes and R-squared values close to zero (Gear Systems: $y=0.0010x+0.5547$, p = 0.981628, R² = 0.000023; Fluid Systems: $y=0.0169x+0.3695$, p = 0.787137, R² = 0.003097). These findings suggest that while model size often correlates with improved performance in many mechanical reasoning domains, some domains may not benefit significantly from increased model scale. This indicates a potential distinction between the underlying mechanisms across different domains of mechanical reasoning.
Additionally, in conjunction with participants’ scores on the Raven’s Advanced Progressive Matrices (APM), we found that that humans’ performance correlates significantly with fluid intelligence (Pearson correlation coefficient $R^{2}=0.24$). This holds for all evaluating dimensions with the exception of System Stability, which exhibited a moderately negative correlation with fluid intelligence (Figure 6). This irregularity likely indicates that tasks in this category rely more on intuitive judgment than on advanced logical reasoning. This static relationship between System Stability and fluid intelligence, however, is not replicated between model performance on said domain and model size. This result potentially indicated a divergence between the underlying process of models and humans in solving System Stability tasks, where models rely more on scalable abilities as opposed to intuitive judgment.
@@ 4 Discussions
Figure 5: The Relationship Between Model Performance and Model Size The six plots on the left illustrate how model size, as measured by numbers of parameter in the models’ neural networks (log scale), affects accuracy across six evaluation dimensions. The plot on the right shows the overall accuracy across tasks.
Here, we constructed MechBench, an assay of 155 cognitive experiments that include six aspects of mechanical reasoning. We investigated VLMs’ ability in reasoning about System Stability, Gear Systems, Pulley Systems, Leverage Principle, Inertia and Motion, and Fluid Mechanics. Despite falling behind humans on all domains, in general they achieved significantly above chance performance, indicating certain levels of mechanical reasoning abilities.
@@
Notably, VLMs’ performance on two dimensions—Gear Systems and Fluid Mechanics—remains close to chance and does not improve with increased model size, serving as a counterexample to the scaling law hypothesis in machine learning (Sutton, 2019; Kaplan et al., 2020). Such a trend indicate that the underlying architecture of these models is yet to be able to support certain mechanisms that are required for a system to have mechanical reasoning abilities regarding Gear System and Fluid Mechanics. Similar trends have previously been reported in VLMs’ performance on level-2 perspective-taking tasks, in which the models are asked to tell how another agent would see the same spatial arrangement in a perspective different from their own (Gao et al., 2024; Piaget & Inhelder, 1969; Moll & Meltzoff, 2011). Interestingly, like mechanical reasoning, said ability is also known to be supported by mental simulation in humans (Hegarty, 2004; Zhao et al., 2016). Together, these findings suggest that VLMs may lack the ability to perform model-based reasoning, possibly highlighting a fundamental limitation in the architecture of the foundational models used in current VLMs.
Figure 6: The Relationship Between Human Performance on MechBench and Human Fluid Intelligence. We further validate our MechBench using correlations with human fluid intelligence, measured via Raven’s Advanced Progressive Matrices (Raven et al., 2000). Interestingly, System Ability seems to be an innate cognitive ability that’s orthogonal to human fluid intelligence.
@@
There is however an important concern with this interpretation: if the limitations of VLMs in mechanical reasoning are indeed due to their inability to perform model-based reasoning, why do they not also under-perform in domains like Pulley Systems, which are also known to rely on such reasoning? Two possible explanations could account for this discrepancy. Foremost, the role of mental simulations in mechanical reasoning is known to be mediated by low-level cognitive abilities, such as intuitions about object relations and motions—often referred to as intuitive physics (McCloskey, 1983a; Kubricht et al., 2017). Intuitive physics involves understanding how objects and substances in the environment are subject to physical laws without relying on formal, abstract knowledge. While intuitive physics may draw on domain-general abilities like model-based reasoning, it also depends on the differential application of key physical principles, such as motion, gravity, and states of matter (Kaiser et al., 1986; Kubricht et al., 2017). Different domains of mechanical reasoning are supported by an intuitive understanding of these physical principles. For instance, reasoning about fluid mechanics heavily relies on understanding liquid properties, which differentiates from that of solid objects (Kawabe et al., 2015). It is possible that models can perform model-based reasoning at a lower level but fail to apply it effectively to certain physical domains, explaining their varied performance across mechanical reasoning tasks.
On the other hand, it may just be that models have learned shortcuts that allow them to bypass model-based reasoning when tackling certain domains. Pulley Systems, for example, are heavily represented in online materials related to physics education. Given that Pulley Systems often involve a small number of components (e.g., fixed and moving pulleys), with limited variety in their arrangements, models may exploit spurious correlations between benchmark questions and similar examples already present in their training data. In other words, the performance of models on certain domains of mechanical reasoning may not reflect genuine, step-by-step reasoning but rather statistical correlations (Bender et al., 2021; Bleeker et al., 2024). This is supported by the observation that very large models, such as GPT-4o, perform particularly well on Pulley System tasks, while smaller models tend to perform much worse.

### v4

@@ Abstract
Mechanical reasoning is a hallmark of human intelligence, underpinning activities from everyday tool use to civil engineering. Endowing machines with this ability is thus a key step toward human-level AI. We evaluate 26 Vision-Language Models (VLMs) on 155 cognitive experiments spanning six domains—system stability, pulley systems, gear systems, the leverage principle, inertia and motion, and Liquid Mechanics. VLMs are showed the weakest on gear systems and Liquid Mechanics. Moreover, performance in these domains does not improve with model size, challenging simple scaling expectations. These results suggest that current attention-based architectures may lack mechanisms needed for mechanical reasoning, particularly those associated with mental simulation.
Keywords: mechanical reasoning; vision-language models; model-based reasoning; intuitive physics
@@ 1 Introduction
Humans are uniquely capable of working with complex mechanical systems, from routine tasks such as assembling furniture to large-scale endeavors such as architectural design and advanced engineering (Harari, 2014). These capabilities rest on the ability to reason about relations and interactions among physical objects—mechanical reasoning (Clark, 2010; Harman, 2011; Vaesen, 2012). While some animals exhibit limited forms of tool use or object interaction (Shumaker et al., 2011), human mechanical reasoning is unparalleled in flexibility, sophistication, and creativity, enabling innovation, problem solving, and adaptation across environments (Allen et al., 2020; Allen, 2021). Mechanical reasoning is therefore a cornerstone of human intelligence and a driver of technological and cultural progress. Endowing artificial systems with this ability is a vital step toward human-level performance in real-world settings. With rapid advances in large language models (LLMs) and their multimodal variants—Vision-Language Models (VLMs)—mechanical reasoning provides a critical lens for assessing current capabilities and limitations.
Although mechanical reasoning is a high-level ability that emerges relatively late in development, it relies on more foundational strategies (Hegarty et al., 1988; Kim & Spelke, 1999; Allen, 2021). Decades of cognitive science research highlight the central role of mental simulation—constructing and manipulating internal models of the world to guide inference (Hegarty, 2004). Interviews and eye-tracking studies show that people dynamically build spatial representations of gears and pulleys to infer behavior while simulating motion (Hegarty, 1992; Lehrer & Schauble, 1998; Kubricht et al., 2017; Rozenblit et al., 2002). By contrast, several studies suggest that the capacity to construct and exploit such visuospatial models remains underdeveloped in state-of-the-art LLMs/VLMs (Hao et al., 2023; Mitchell & Krakauer, 2023; Goddu et al., 2024; Gao et al., 2024). Because mental simulation requires model-based reasoning in the visuospatial domain, a thorough evaluation of VLMs’ mechanical reasoning can inform this debate. In this work, we leveraged the MechBench from the CoreCognition benchmark (Li et al., 2024b), which comprises 155 single-image tasks adapted from cognitive psychology that spans six core domains of mechanical reasoning: System Stability, Pulley Systems, Gear Systems, Leverage Principle, Inertia and Motion, and Liquid Mechanics. Conducting the first systematic evaluation of 26 open- and closed-source VLMs against a human baseline, we reveal consistent human superiority, notable domain-specific performance gaps, and weak scaling-law effects in certain domains such as gear systems and Liquid Mechanics. Finally, we provide a cognitive-science-informed analysis linking VLM performance limitations to deficits in model-based reasoning, offering insights into architectural constraints and shortcut exploitation in current models.
@@ 2 Methods
2.1 Experiment Design
Figure 1: Sample tasks across the six MechBench domains. For each domain, GPT-4o answered correctly on the left and failed on the right.
Mechanical reasoning has been widely explored in cognitive psychology and educational research. This study uses single-image adaptations of classic cognitive tasks from the literature to assess mechanical reasoning, focusing on six key domains: system stability, pulley systems, gear systems, leverage principle, inertia and motion, and liquid mechanics. Example cases on each domains are presented in Figure 1. This categorization is intentionally designed to encompass the full range of mechanical reasoning, ensuring each category represents a distinct but fundamental physical understanding. Below, we provide explanations for each domain.
@@ 2.1 Experiment Design
Liquid Mechanics
Liquid Mechanics involves understanding the behavior of liquids under various conditions, such as flow, external pressures, and volume changes. Although grounded in the intuitive understanding of fluid dynamics emerged very early in humans’ cognitive development (Hespos et al., 2016), reasoning about liquid behaviors in mechanical systems require simultaneous consideration of geometry, force, and dynamics. These tasks represent a highly important domain of mechanical reasoning concerning specifically about liquid motions as opposed to solid objects alone.
2.2 Examined Vision Language Models
Recent advances in multi-modal learning have been driven by the unified modeling of visual and textual modalities using transformers (Li et al., 2019; Xu et al., 2023; Tan & Bansal, 2019; Alayrac et al., 2022; Radford et al., 2021). With the rise of LLMs, state-of-the-art (SOTA) multi-modal LLMs (MLLMs) (Liu et al., 2024b; Li et al., 2023b) adopt open-source LLMs (Touvron et al., 2023; Peng et al., 2023; Jiang et al., 2023) and align visual features to the LLM embedding space (Li et al., 2023a; Fu et al., 2023; Wu & Xie, 2024; Xu et al., 2024; Shao et al., 2024; Li et al., 2022; 2025; Brown et al., 2020; Achiam et al., 2023; Bai et al., 2023a; Jaech et al., 2024; Zhang et al., 2025; 2024). Progressively, MLLMs have demonstrated competitive performance in complex tasks involving high-level perception and reasoning (Li et al., 2024a; Gemini, 2023; Fu et al., 2023; OpenAI, 2023), such as spatial reasoning (Chen et al., 2024a; Cai et al., 2024), character recognition (Mori et al., 1999), scene understanding (Cordts et al., 2016; Wang et al., 2023; Li et al., 2023c; Chen et al., 2024c; b), action recognition (Jhuang et al., 2013; Herath et al., 2017) and prediction (Lan et al., 2014; Kong & Fu, 2022), reaching near-human performance.
@@ 2.2 Examined Vision Language Models
    Open-source VLMs with Multi-Image Reasoning: Includes models with different sizes and other variants such as CogVLM Series(Hong et al., 2024), Qwen series(Qwen-VL (Bai et al., 2023b), Qwen-2 (Wang et al., 2024)), and Blip2 (Li et al., 2023b), LLaVA-Next (Liu et al., 2024a) , which are capable of reasoning over interleaved multiple images and texts.
@@ 2.3 Human Baseline
We recruited a total of 9 participants, all of whom were college students proficient in English. Participants were instructed to skip any question that was ambiguously phrased or too complex to answer within 90 seconds. As a sanity check, a question was marked as failed if less than 80% of the participants could answer within the time frame; otherwise, we modified the question, and new annotators completed the revised version.
@@ 3 Results
Our study reveals a significant disparity between human and model performance across multiple evaluation dimensions. As shown in Figure 2 and 3, humans consistently outperform models both in overall accuracy and in task-specific domains. These results highlight the limitations of current VLMs in replicating human-like reasoning in mechanical and intuitive physics tasks. Among the evaluating domains, Pulley Systems exhibits the largest performance gap, with human accuracy nearing 90%, compared to model accuracy averaging around 50%.
@@
Figure 3: Overall and Domain-Wise Accuracy: Humans vs. Models. A. For overall accuracy across tasks, human participants outperform models significantly (p < 0.0001). B. Human participants consistently outperform models in each domain (all categories p < 0.001).
Figure 4: Comparison of Tukey HSD P-values for Model and Human Performance Across Different domains.
3.1 Human vs. Model Performance Across Each Domain
The ANOVA test ($F=11.8111,p<0.0001$) revealed significant differences in human performance across the six categories. Tukey HSD post-hoc analysis (visualized in the heatmap) shows that the performance of humans handling Pulley Systems is significantly different from the other five categories. The performance of humans handling the pulley system is significantly better than the other five categories. In contrast, ANOVA results for model performance ($F=1.6809,p=0.1430$) indicated no significant differences across task categories. Tukey HSD analysis further reveals that models exhibited similar accuracy across all categories with the exception of Liquid Mechanics, which deviates significantly with other domains. Together, as shown in Figure 3, the distribution of human performance is narrower, indicating more consistent accuracy, whereas the models display a wider range of variability. See Appendix A for detailed analysis on representative model response and reasoning for each domain.
@@ 3.2 Relationship Between Model Performance and Model Size
A widely held belief in the machine learning community is that an increase in a model’s scale, measured by the number of parameters, leads to systematic improvements in its reasoning abilities (Sutton, 2019; Kaplan et al., 2020), a concept known as the scaling law. However, this assumption is an empirical observation without theoretical proof (Luo et al., 2025). To evaluate whether the scaling law applies to mechanical reasoning, we further examined the relationship between model performance on mechanical reasoning tasks and model size, as measured by the number of parameters (Figure 5). For overall accuracy across task, regression analysis yielded a formula $y=0.0821x+0.4318$ (p = 0.0053, R² = 0.2917). Models with parameter counts exceeding 10 billion, such as GPT-4o High and GPT4V-TURBO-HIGH, distinctively outperformed smaller models. In most of the domains (System Stability, Pulley Systems, Leverage Principle, Inertia and Motion, Liquid Mechanics), accuracy increases with larger model sizes. However, Gear Systems and Liquid Mechanics show very weak scaling with slopes and R-squared values close to zero (Gear Systems: $y=0.0010x+0.5547$, p = 0.981628, R² = 0.000023; Liquid Mechanics: $y=0.0169x+0.3695$, p = 0.787137, R² = 0.003097). While model size often correlates with improved performance in many mechanical reasoning domains, some domains may not benefit significantly from increased model scale, indicating a potential distinction between the underlying mechanisms across different domains of mechanical reasoning.
Figure 5: Relationship between model performance and model size. The six panels on the left show how model size—measured by parameter count (log scale)—relates to accuracy across the six evaluation domains. The panel on the right reports overall accuracy across all tasks.
@@ 4 Discussions
In this paper, we examine domain-specific patterns in VLM performance on mechanical reasoning tasks, showing that accuracy lags well behind human levels across all domains. While larger models tend to achieve higher overall accuracy, the strength of this relationship varies markedly by domain. Notably, VLMs’ performance on Gear Systems and Liquid Mechanics remains close to chance and does not improve with increased model size, serving as a counterexample to the scaling law hypothesis in machine learning (Sutton, 2019; Kaplan et al., 2020). Such a trend indicate that the underlying architecture of these models is yet to be able to support certain mechanisms that are required for a system to have mechanical reasoning abilities regarding Gear System and Liquid Mechanics. Similar trends have previously been reported in VLMs’ performance on level-2 perspective-taking tasks, in which the models are asked to tell how another agent would see the same spatial arrangement in a perspective different from their own (Gao et al., 2024; Piaget & Inhelder, 1969; Moll & Meltzoff, 2011). Interestingly, like mechanical reasoning, such ability is also known to require mental simulation in humans, which is the internal construction and manipulation of dynamic models of the world to predict or explain outcomes (Hegarty, 2004; Zhao et al., 2016). In cognitive science, mental simulation is often considered a concrete instantiation of model-based reasoning, where inferences are produced by running an internal model to project likely outcomes, rather than relying solely on pre-learned patterns or correlations (Johnson-Laird, 1983; LeCun, 2022; Balaban & Ullman, 2025). Together, these findings suggest that VLMs may lack the capacity for such simulation-driven, model-based reasoning, possibly highlighting a fundamental limitation in the architecture of current foundational models (Mitchell & Krakauer, 2023; Goddu et al., 2024; Li et al., 2024b).
@@
There is, however, an important concern with this interpretation: if VLMs’ limitations in mechanical reasoning stem from an inability to perform model-based reasoning, why do they not also underperform on domains like Pulley Systems, which likewise rely on such reasoning? Two explanations may account for this discrepancy. First, the role of mental simulation is mediated by lower-level cognitive abilities often termed intuitive physics—expectations about object relations and motion (McCloskey, 1983a; Kubricht et al., 2017). Intuitive physics draws on, but is not reducible to, model-based reasoning; it depends on the domain-specific application of physical primitives such as motion, gravity, and states of matter (Kaiser et al., 1986; Kubricht et al., 2017). Different domains recruit different primitives. For example, Liquid Mechanics relies on priors about liquids that differ from those governing solids (Kawabe et al., 2015). Models may therefore exhibit some model-based reasoning yet fail to apply the appropriate primitives in certain domains, yielding a gap in performance.
Second, models may exploit shortcuts that bypass genuine simulation. Pulley systems are heavily represented in educational materials and typically involve few components (e.g., fixed and movable pulleys) with limited combinatorics, enabling spurious correlations between the benchmarked domain and patterns present in training data (Bender et al., 2021; Bleeker et al., 2024). Apparent competence may thus reflect pattern matching rather than step-by-step reasoning. Consistent with this account, very large models (e.g., GPT-4o) perform particularly well on pulley tasks, whereas smaller models lag.